At 9:00 a.m. a truck leaves the truck yard and travels west at a rate of 45 mi/hr. Two hours later, a second truck leaves along the same route, traveling at 55 mi/hr. When will the second truck catch up to the first?

Thank you!!!

Answer by rajagopalan(148)

start time of truck 1 = 9
start time of truck 2 = 11
Position of truck 1 at 11 am = 45*2 = 90 km apart
speed of truck 2 =55
speed of truck 1 =45
Apprach velocity = 55-45 = 10 mph
So to cover 90 miles time needed = 90/10=9 hours
Second truck will catch up / overtake 1st truck at 11+9 = 20 hours
or (08:00 pm)on the same day.
Cheers!